Could you please solve this system by the addition method, indicating if the system is independent, inconsistent, or dependent.
I really appreciate your help. Thank you very much.
-3x + y = 3
2x - 3y = 5

Answer by oscargut(2103)

-3x + y = 3 (x3)
2x - 3y = 5
then
-9x+3y=9
2x - 3y = 5
then doing the addition
-7x=14
then
x=-2
then
2(-2)-3y=5
then
-3y=9
then
y=-3
Solution is x=-2 and y=-3 (unique)
so the system is independent